OpenStudy (anonymous):

HELP FAST I WILL FAN AND MEDAL the probability of having a winning raffle ticket is 20%. if you brought 50 tickets, how many winning tickets should you expect to have? A. 3 tickets B. 8 tickets C. 10 tickets D. 20 tickets

OpenStudy (pinksapphire):

\[\frac{ 20 }{ 100 }=\frac{ x }{ 50}\]

OpenStudy (pinksapphire):

The correct answer would be C.

OpenStudy (pinksapphire):

Because 100/50=2 so..20/x=2

OpenStudy (anonymous):

1/50

OpenStudy (jim766):

20% of the tickets should be winners. change 20% to a decimal (divide it by 100) (20/100) * 50 =

OpenStudy (anonymous):

i got C

OpenStudy (jim766):

yup, same as @PinkSapphire ,